# KACQ-DCNN: Uncertainty-Aware Interpretable Kolmogorov–Arnold Classical–Quantum Dual-Channel Neural Network for Heart Disease Detection

### v3

@@ Abstract
QML is an emerging field that combines quantum computing with traditional ML techniques, offering the potential to enhance computational power and improve diagnostic accuracy. Although still in its early stages, QML has shown potential for medical applications. Magallanes et al. [42] demonstrated the superiority of a hybrid classical–quantum model over CMLs in detecting stenosis from X-ray coronary angiography, significantly improving accuracy, recall, and F1 scores. Gupta et al. [43] compared QML techniques with DL models for diabetes prognosis using the PIMA Indian Diabetes dataset. Although the DL model achieved higher accuracy (95%), the QML model still performed well, showing promise in handling large datasets. Maheshwari et al. [44] explored the computational efficiency of quantum algorithms compared with classical algorithms in diabetes classification and reported that quantum approaches improved computational speed by up to 55 times. This suggests that although QML may not yet consistently outperform traditional models in terms of accuracy, its ability to handle vast datasets efficiently positions it as a valuable tool for medical prediction. Kumar et al. [23] compared traditional ML with QML algorithms and reported that the QML algorithm achieved an accuracy of 89% in heart failure detection, with notable improvements in F1-score and recall. These findings suggest that although QML is still being developed, its advantages in handling large, complex datasets and improving computational efficiency are promising. QML has also gained traction; Kavitha and Kaulgud [24] presented a quantum K-means algorithm that attained 96.40% accuracy, marking significant progress in applying quantum computing to heart disease diagnostics. While current applications of QML are limited, the field is rapidly evolving, and future advancements may allow quantum models to surpass classical methods entirely.
@@
KANs represent the latest innovative approach to neural network architectures, offering enhanced accuracy with fewer parameters, dynamic adaptability, and architectural flexibility in medical diagnostics. Li et al. [45] developed the U-KAN model, a hybrid approach that combines KANs with U-Net for medical image segmentation tasks. This model reduces computational costs and improves segmentation performance, making it highly suitable for applications such as MRI analysis. In addition, KANs have been used for seizure detection based on EEG data. Contreras et al. [46] developed the KAN-EEG model, which demonstrated superior generalizability across datasets from different regions, reducing overfitting and improving diagnostic accuracy. The flexibility and adaptability of KANs in medical diagnostics make them valuable, particularly for tasks requiring high interpretability and computational efficiency.
@@
ML has proven to be a valuable tool for heart disease prediction, with many models showing high accuracy and reliability. Although there are multiple databases related to heart disease, most studies have been conducted via the Cleveland database, and a few have used Statlog, demonstrating the need for more diverse and aggregated datasets. Conventional neural networks typically demand a higher number of parameters and extensive computational resources, posing a significant challenge in environments with limited resources. Traditional ML algorithms such as RF and SVM have demonstrated robustness in clinical settings, while hybrid and DL models offer enhanced accuracy for complex datasets. However, previous research focused only on a single type of dataset at a time, which might contain potential bias, lack of generalizability, or overfitting because of being short, highly categorical, and imbalanced. While previous studies explored QML and CQML in heart disease prediction [2, 24, 23], showing their promise for improving computational efficiency and diagnostic performance, they have yet to outperform classical methods consistently. KANs, as substitutes for MLP, offer a novel approach to improving the accuracy with less computational complexity and flexibility of neural network architectures in medical diagnostics, which remains greatly unexplored. Integrating KANs with CQML techniques will likely lead to more precise, efficient, and interpretable models for heart disease prediction.
Methods
@@
Figure 1 illustrates the comprehensive methodology employed for developing and evaluating the proposed hybrid KACQ-DCNN framework, which is discussed in subsequent subsections. It begins with the data acquisition phase, in which heart disease datasets from multiple sources are integrated. These datasets undergo preprocessing steps such as handling missing values, categorical encoding, data normalization, outlier detection, balancing, adding feature interaction terms, and splitting the dataset. In the model training phase, CML classifiers were initialized with GridSearchCV, and cross-validation was performed to select the best-performing model based on the evaluation metrics. Once the best model was identified, it predicted the labels on the test set and generated a classification report to evaluate its performance. Using the class probabilities, we combined them with the original feature sets to create an enhanced feature set for both training and test sets. The KACQ-DCNN architecture incorporates classical and dressed quantum models as a part of the hybrid dual-channel model. The engineered features are passed as inputs through each channel through which the data are processed in parallel. The output from the dual channel was concatenated into a final layer, followed by a 2-neuron output layer with sigmoid activation for binary classification. The evaluation phase computes each class’s classification metrics, predicted labels, and probabilities. The model includes uncertainty quantification and applies post-hoc explainability techniques, ensuring that it is interpretable and reliable for clinical use.
Dataset description
@@
Notably, asymptomatic (ASY) is most common in heart disease patients, whereas atypical angina (ATA) predominates in non-heart disease subjects, as shown in Figure LABEL:fig:chest_pain_type. Similarly, Figure LABEL:fig:peak_excersize shows a higher prevalence of ‘Flat’ slopes observed in heart disease cases, suggesting potential diagnostic value. Figure 4 displays noteworthy patterns, including vs. showing a positive correlation, levels displaying a bimodal distribution in heart disease patients, negatively correlated with age, and values generally greater in heart disease cases. This figure identifies complex, multivariable relationships that may inform risk assessment and diagnostic strategies for heart disease.
Experimental setup
@@
Classical conventional ML models
@@
Quantum neural networks
@@
TTN
TTN [47, 48] provides a hierarchical structure for modeling complex quantum states and is particularly effective for capturing correlations at multiple scales. Unlike simpler tensor network models such as MPS, TTN organizes tensors in a tree-like structure. This provides a more detailed picture of how subsystems interact and become entangled.
@@
Proposed KACQ-DCNN
@@
Architecture of quantum circuit
@@
Next, measurement operator is defined on the basis of the desired output. Heart disease prediction can be modeled as a projection operator onto a subspace defined by a binary outcome (e.g., presence or absence of heart disease). Mathematically, this operator can be represented as:
@@
Integrating classical and dressed quantum circuit
Integrating the BiLSTMKANnet and QDenseKANnet models is designed as a dual-channel architecture that leverages both classical DL and quantum-enhanced computations. The architecture begins by processing the input data separately through two channels: classical BiLSTMKANnet and hybrid quantum-classical QDenseKANnet. Both models accept input tensors of shape , meaning samples with 12 features, each of them represented as a single value. and 1 can be interpreted as the feature dimension, commonly used in neural networks to handle input with multiple features in a consistent shape. However, due to their strengths, the data undergo distinct transformations within each network.
@@
Training and tuning process
@@
For stratified 10-fold cross-validation, all models went through 10 rounds of experiments, each comprising a split, training, and testing phase. The dataset was divided into 10 equal folds in each round, with 9 folds used for training and 1 fold for testing. This process was repeated 10 times to ensure that each fold was used for testing exactly once. The reported ‘’ values in Table 3 and 5 represent the mean and standard deviation of the performance metrics across these 10 rounds, providing a robust measure of both the average performance and the variability for each model.
@@
Similarly, DenseKAN layer units were searched for between 32 and 256. Hyperband was used with a maximum of 10 epochs per trial, and early stopping with patience for 3 epochs was applied to prevent overfitting. This process yielded the best configuration by maximizing the validation accuracy across 50 epochs, and the optimal set of hyperparameters was extracted from one trial. For QCKANnet, the tuner explored filters for Conv1DKAN layers in the range of 32-128 in step 32. The Keras tuner used the same settings as in BiLSTMKANnet to maximize validation accuracy while adhering to the early stopping criterion. Finally, QDenseKANnet was used to evaluate the first, second, and fourth DenseKAN layers from 16 to 256 units in step 32. The best configuration was selected on the basis of validation accuracy, and the hyperband facilitated rapid convergence through efficient resource allocation.
Results and Discussion
Model performance
Evaluation metrics
@@
Macro average F1-score (maF1): The F1-score is the harmonic mean of precision and recall. It balances precision and recall, making it useful when the costs of false positives and negatives are high. maF1 averages the F1 scores across all classes:
@@
Model Category | Models | Evaluation Metrics | Training Time (s) | Inference Time (s) | ||||||
maP | maR | maF1 | Accuracy | ROC-AUC | MCC | Kappa | ||||
@@
Author(s) | Dataset | Model | Evaluation Metrics | |||
maP | maR | maF1 | Accuracy | |||
Singh and Kumar[11] | Cleveland dataset | KNN Classifier (k=3) | 87.76% | 72.88% | 79.54% | 87.00% |
Kavitha et al.[15] | Cleveland dataset | Hybrid Classifier (Decision Tree + Random Forest | - | - | - | 88.00% |
Mohan et al.[17] | Cleveland dataset | HRFLM (proposed) | 90.10% | 92.80% | 90.00% | 88.40% |
Sharma et al.[10] | Cleveland dataset | Random Forest | 99.00% | 99.00% | - | - |
Saboor et al.[9] | Cleveland dataset | SVM Classifier | 94.00% | 94.00% | 94.00% | 96.72% |
Rani et al.[8] | Cleveland dataset | Random Forrest | 88.46% | 84.14% | 86.25% | 86.60% |
Shah et al.[49] | Cleveland dataset | KNN Classifier (k=7) | - | - | - | 90.78% |
Li et al.[5] | Cleveland dataset | FCMIM (The feature selector algorithm which led to better results) | - | 92.50% | - | 92.00% |
Doppala et al.[50] | Cleveland dataset | Genetic Algorithm + Radial Basis Function | 95.00% | 96.00% | 95.00% | 94.00% |
Kavitha and Kaulgud[24] | Cleveland dataset | Quantum K-means | 96.80% | 95.00% | 96.40% | 96.40% |
Kumar et al.[23] | Cleveland dataset | QRFC | 89.00% | 93.00% | 88.00% | 89.00% |
Babu et al.[2] | Cleveland dataset | QuEML | 93.00% | 94.00% | 93.00% | 93.00% |
Mir et al.[16] | Heart Statlog dataset | Random Forest + Adaboost | 99.40% | 99.69% | 99.55% | 99.48% |
Mir et al.[16] | Heart disease UCI dataset | Random Forest + Adaboost | 92.23% | 96.75% | 94.44% | 93.90% |
Karadeniz, Tokdemir, and Maraş[51] | Heart Statlog dataset | Shrunk Covariance classifier | 85.50% | 89.80% | 87.60% | 88.80% |
Anuradha and David[52] | Heart Statlog dataset | Majority Voting Ensemble | 87.00% | 87.00% | 87.00% | 87.04% |
Patro, Nayak, and Padhy[53] | Heart disease UCI dataset | Bayesian Optimized Support Vector Machine | 100.00% | 80.00% | - | 93.30% |
Performance benchmarking
@@
In contrast to the existing literature, mentioned in Table 4, where models were trained on individual heart disease datasets, our approach is novel in that we merged multiple datasets, including the Statlog and UCI dataset, to show generalizability and robustness of KACQ-DCNN. This is the first attempt in the literature to train a model using such a merged heart disease dataset, offering a more holistic view across different datasets. It’s important to highlight that the Cleveland dataset is a small subset of the UCI dataset (Cleveland, Hungarian, Switzerland, and Long Beach, Virginia), on which several studies trained models individually. Our KAN-enhanced method allows for broader model generalization, potentially capturing patterns across datasets, which could improve robustness in real-world applications. Despite this complexity, our proposed KACQ-DCNN achieved 92% across all evaluation metrics, indicating strong performance even when dealing with the diversity of combined datasets. In the literature, dataset-specific models were used to achieve high accuracy using only one dataset. For example, Mir et al. [16] got 99.48% accuracy via RF + Adaboost on the Heart Statlog dataset. However, KACQ-DCNN also achieved superior state-of-the-art performance on Cleveland and Statlog datasets, both sharing an equal maF1 of 96% and the accuracy of 96.23% and 96.30%, respectively.
@@
Over 16 epochs, KACQ-DCNN initially showed consistent improvements in terms of loss and accuracy. The first epoch resulted in a loss of 0.7087, with a validation accuracy of 73.55%. By the third epoch, validation accuracy reached 88.77%, and by the seventh epoch, it peaked at 92.03%. The accuracy subsequently fluctuated without significant improvements. The learning rate was reduced in epoch 12 to enhance convergence from to . However, validation accuracy could not surpass the 92.03% peak, and training was halted at epoch 16 owing to the early stopping mechanism. As shown in Figure LABEL:fig:loss, the training loss steadily decreased across all epochs, indicating effective learning. However, the validation loss plateaued around epoch 7, with minor fluctuations suggesting that the model was near its generalization capacity. Despite the learning rate reduction at epoch 12, validation loss showed no substantial gains, confirming the onset of overfitting. As depicted by Figure LABEL:fig:cm, KACQ-DCNN correctly predicts 40.22% of normal cases and 51.81% of HF cases, indicating good performance in detecting HFs. However, 4.35% of normal cases were misclassified as HF, whereas 3.62% of HF cases were incorrectly labeled as normal.
@@
Figures LABEL:fig:roc_curve and LABEL:fig:pr_curve illustrate the performance of KACQ-DCNN via the ROC and precision-recall curves, which are critical metrics for evaluating binary classification tasks. The ROC curve demonstrates a high TP rate against the FP rate, indicating that the model effectively distinguishes between the positive and negative classes. An AUC score of 99.73% on the training set and 94.77% on the test set suggested excellent classification ability, with minimal risk of overfitting. Similarly, the precision-recall curve reveals the model’s ability to maintain high precision and recall rates, achieving values of 99.78% in training and 94.89% in testing. This finding indicates that the model is accurate and robust across both datasets, demonstrating its effectiveness in predicting the target classes.
@@
Ablation study analysis

### v4

@@ 2.2 QML in Medical Prediction
QML is an emerging field that combines quantum computing with traditional ML techniques, offering the potential to enhance computational power and improve diagnostic accuracy. Although still in its early stages, QML has shown potential for medical applications. Magallanes et al. (Ovalle-Magallanes et al.,, 2022) demonstrated the superiority of a hybrid classical–quantum model over CMLs in detecting stenosis from X-ray coronary angiography, significantly improving accuracy, recall, and F1 scores. Gupta et al. (Gupta et al.,, 2022) compared QML techniques with DL models for diabetes prognosis using the PIMA Indian Diabetes dataset. Although the DL model achieved higher accuracy (95%), the QML model still performed well, showing promise in handling large datasets. Maheshwari et al. (Maheshwari et al.,, 2020) explored the computational efficiency of quantum algorithms compared with classical algorithms in diabetes classification and reported that quantum approaches improved computational speed by up to 55 times. This suggests that although QML may not yet consistently outperform traditional models in terms of accuracy, its ability to handle vast datasets efficiently positions it as a valuable tool for medical prediction. Kumar et al. (Kumar et al.,, 2021) compared traditional ML with QML algorithms and reported that the QML algorithm achieved an accuracy of 89% in heart failure detection, with notable improvements in F1-score and recall. These findings suggest that although QML is still being developed, its advantages in handling large, complex datasets and improving computational efficiency are promising. QML has also gained traction; Kavitha and Kaulgud (Kavitha and Kaulgud,, 2023) presented a quantum K-means algorithm that attained 96.40% accuracy, marking significant progress in applying quantum computing to heart disease diagnostics. While current applications of QML are limited, the field is rapidly evolving, and future advancements may allow quantum models to surpass classical methods entirely.
Study | Method(s) | Dataset | Accuracy | Highlights |
Ovalle-Magallanes et al., (2022) | CQML (Quantum-Classical Hybrid) | Coronary Angiography | 91.80% | Outperformed CML baselines |
Gupta et al., (2022) | QML vs. DL | PIMA Diabetes | QML: 86% | QML promising for large data |
Maheshwari et al., (2020) | QML vs. CML | Diabetes | QML: 69% | QML was 55x faster |
Kumar et al., (2021) | QML vs. Classical ML | Heart Failure | QML: 89% | Better F1 and Recall scores |
Babu et al., (2024) | QML vs. CML | Cleveland | QuEML: 93% | Outperformed CML benchmarks |
Kavitha and Kaulgud, (2023) | Quantum K-means | Cleveland | 96.40% | Major step for QML diagnostics |
Contreras et al., (2024) | KAN-EEG | EEG Seizure Data | – | Generalizable and reduced overfitting |
Li et al., (2024) | Seg. U-KAN (KAN + U-Net) | Medical Imaging | – | Accurate and computationally efficient |
2.3 KANs in Medical AI
@@
KANs represent the latest innovative approach to neural network architectures, offering enhanced accuracy with fewer parameters, dynamic adaptability, and architectural flexibility in medical diagnostics. Li et al. (Li et al.,, 2024) developed the U-KAN model, a hybrid approach that combines KANs with U-Net for medical image segmentation tasks. This model reduces computational costs and improves segmentation performance, making it highly suitable for applications such as MRI analysis. In addition, KANs have been used for seizure detection based on EEG data. Contreras et al. (Contreras et al.,, 2024) developed the KAN-EEG model, which demonstrated superior generalizability across datasets from different regions, reducing overfitting and improving diagnostic accuracy. The flexibility and adaptability of KANs in medical diagnostics make them valuable, particularly for tasks requiring high interpretability and computational efficiency.
@@
ML has proven to be a valuable tool for heart disease prediction, with many models showing high accuracy and reliability. Although there are multiple databases related to heart disease, most studies have been conducted via the Cleveland database, and a few have used Statlog, demonstrating the need for more diverse and aggregated datasets. Conventional neural networks typically demand a higher number of parameters and extensive computational resources, posing a significant challenge in environments with limited resources. Traditional ML algorithms such as RF and SVM have demonstrated robustness in clinical settings, while hybrid and DL models offer enhanced accuracy for complex datasets. However, previous research focused only on a single type of dataset at a time, which might contain potential bias, lack of generalizability, or overfitting because of being short, highly categorical, and imbalanced. While previous studies explored QML and CQML in heart disease prediction (Babu et al.,, 2024; Kavitha and Kaulgud,, 2023; Kumar et al.,, 2021), showing their promise for improving computational efficiency and diagnostic performance, they have yet to outperform classical methods consistently. KANs, as substitutes for MLP, offer a novel approach to improving the accuracy with less computational complexity and flexibility of neural network architectures in medical diagnostics, which remains greatly unexplored. Integrating KANs with CQML techniques will likely lead to more precise, efficient, and interpretable models for heart disease prediction.
3 Methods
@@
Figure 1 presents the comprehensive methodology for developing and evaluating our proposed KACQ-DCNN framework. The process initiates with data acquisition, where multiple heart disease datasets from diverse sources are integrated (see Section 3.1). These datasets then undergo extensive preprocessing and exploratory data analysis, including handling missing values, categorical encoding, normalization, outlier detection and management, data balancing through SMOTE, generation of feature interaction terms, and dataset splitting into training and test sets (detailed in Section 3.7). In the classical feature engineering phase, various classical machine learning (CML) classifiers are initialized with GridSearchCV and evaluated via cross-validation to identify the optimal model based on multiple performance metrics. The selected model not only predicts labels on the test set but also outputs class probabilities and a detailed classification report. Additionally, model-derived information, such as leaf indices or probability scores, is combined with the original features to form an enriched feature set for both training and testing data. At the heart of our framework lies the KACQ-DCNN architecture (explained in Section 3.8), a hybrid dual-channel model integrating both classical (BiLSTMKANnet) and dressed quantum (QDenseKANnet) components. These engineered features are processed in parallel through each channel, with their outputs concatenated and passed to a final dense layer followed by a 2-neuron sigmoid-activated output layer for binary classification. The subsequent evaluation phase computes comprehensive classification metrics, predicted labels, and class probabilities (Section 4.1.1). Finally, to ensure clinical interpretability and reliability, a post-hoc analysis phase is conducted, involving uncertainty quantification through conformal prediction (Section 4.3) and model explainability using techniques such as SHAP and LIME (Section 4.2).
3.1 Dataset description
@@
Notably, asymptomatic (ASY) is most common in heart disease patients, whereas atypical angina (ATA) predominates in non-heart disease subjects, as shown in Figure 3(a). Similarly, Figure 3(b) shows a higher prevalence of ‘Flat’ slopes observed in heart disease cases, suggesting potential diagnostic value. Figure 4 displays noteworthy patterns, including vs. showing a positive correlation, levels displaying a bimodal distribution in heart disease patients, negatively correlated with age, and values generally greater in heart disease cases. This figure identifies complex, multivariable relationships that may inform risk assessment and diagnostic strategies for heart disease.
3.2 Experimental setup
@@ 3.3 Data preprocessing
3.4 Classical conventional ML models
@@
3.5 Quantum neural networks
@@ 3.5.2 MPS
3.5.3 TTN
TTN (Schollwöck,, 2005; Evenbly and Vidal,, 2009) provides a hierarchical structure for modeling complex quantum states and is particularly effective for capturing correlations at multiple scales. Unlike simpler tensor network models such as MPS, TTN organizes tensors in a tree-like structure. This provides a more detailed picture of how subsystems interact and become entangled.
@@ 3.6.3 QDenseKANnet
3.7 Proposed KACQ-DCNN
@@
3.7.1 Architecture of quantum circuit
@@
Next, the measurement operator is defined based on the desired output. Heart disease prediction can be modeled as a projection operator onto a subspace defined by a binary outcome (e.g., presence or absence of heart disease). Mathematically, this operator can be represented as:
@@ 3.7.2 Architecture of KAN
3.7.3 Integrating classical and dressed quantum circuit
Integrating the BiLSTMKANnet and QDenseKANnet models is designed as a dual-channel architecture that leverages both classical DL and quantum-enhanced computations. The architecture begins by processing the input data separately through two channels: classical BiLSTMKANnet and hybrid quantum-classical QDenseKANnet. Both models accept input tensors of shape , meaning samples with 12 features, each of them represented as a single value, and 1 can be interpreted as the feature dimension, commonly used in neural networks to handle input with multiple features in a consistent shape. However, due to their strengths, the data undergo distinct transformations within each network.
@@ 3.7.3 Integrating classical and dressed quantum circuit
3.8 Training and tuning process
@@
For stratified 10-fold cross-validation, all models went through 10 rounds of experiments, each comprising a split, training, and testing phase. The dataset was divided into 10 equal folds in each round, with 9 folds used for training and 1 fold for testing. This process was repeated 10 times to ensure that each fold was used for testing exactly once. The reported ‘’ values in Table 5 and 7 represent the mean and standard deviation of the performance metrics across these 10 rounds, providing a robust measure of both the average performance and the variability for each model.
@@
Similarly, DenseKAN layer units were searched for between 32 and 256. Hyperband was used with a maximum of 10 epochs per trial, and early stopping with patience for 3 epochs was applied to prevent overfitting. This process yielded the best configuration by maximizing the validation accuracy across 50 epochs, and the optimal set of hyperparameters was extracted from one trial. For QCKANnet, the tuner explored filters for Conv1DKAN layers in the range of 32-128 in steps of 32. The Keras tuner used the same settings as in BiLSTMKANnet to maximize validation accuracy while adhering to the early stopping criterion. Finally, QDenseKANnet was used to evaluate the first, second, and fourth DenseKAN layers from 16 to 256 units in step 32. The best configuration was selected on the basis of validation accuracy, and the hyperband facilitated rapid convergence through efficient resource allocation.
4 Results and Discussion
4.1 Model performance
4.1.1 Evaluation metrics
@@
Macro average F1-score (maF1): The F1-score is the harmonic mean of precision and recall. It balances precision and recall, making it useful when the costs of FPs and FNs are high. maF1 averages the F1 scores across all classes:
@@
Model Category | Models | Evaluation Metrics | Training Time (s) () | Inference Time (s) () | ||||||
maP () | maR () | maF1 () | Accuracy () | ROC-AUC () | MCC () | Kappa () | ||||
@@
4.1.2 Performance benchmarking
@@
Author(s) | Dataset | Model | Evaluation Metrics | |||
maP | maR | maF1 | Accuracy | |||
Singh and Kumar (Singh and Kumar,, 2020) | Cleveland dataset | KNN Classifier (k=3) | 87.76% | 72.88% | 79.54% | 87.00% |
Kavitha et al. (Kavitha et al.,, 2021) | Cleveland dataset | Hybrid Classifier (Decision Tree + Random Forest | - | - | - | 88.00% |
Mohan et al. (Mohan et al.,, 2019) | Cleveland dataset | HRFLM (proposed) | 90.10% | 92.80% | 90.00% | 88.40% |
Sharma et al. (Sharma et al.,, 2020) | Cleveland dataset | Random Forest | 99.00% | 99.00% | - | - |
Saboor et al. (Saboor et al.,, 2022) | Cleveland dataset | SVM Classifier | 94.00% | 94.00% | 94.00% | 96.72% |
Rani et al. (Rani et al.,, 2021) | Cleveland dataset | Random Forrest | 88.46% | 84.14% | 86.25% | 86.60% |
Shah et al. (Shah et al.,, 2020) | Cleveland dataset | KNN Classifier (k=7) | - | - | - | 90.78% |
Li et al. (Li et al.,, 2020) | Cleveland dataset | FCMIM (The feature selector algorithm which led to better results) | - | 92.50% | - | 92.00% |
Doppala et al. (Doppala et al.,, 2023) | Cleveland dataset | Genetic Algorithm + Radial Basis Function | 95.00% | 96.00% | 95.00% | 94.00% |
Kavitha and Kaulgud (Kavitha and Kaulgud,, 2023) | Cleveland dataset | Quantum K-means | 96.80% | 95.00% | 96.40% | 96.40% |
Kumar et al. (Kumar et al.,, 2021) | Cleveland dataset | QRFC | 89.00% | 93.00% | 88.00% | 89.00% |
Babu et al. (Babu et al.,, 2024) | Cleveland dataset | QuEML | 93.00% | 94.00% | 93.00% | 93.00% |
Mir et al. (Mir et al.,, 2024) | Heart Statlog dataset | Random Forest + Adaboost | 99.40% | 99.69% | 99.55% | 99.48% |
Mir et al. (Mir et al.,, 2024) | Heart disease UCI dataset | Random Forest + Adaboost | 92.23% | 96.75% | 94.44% | 93.90% |
Karadeniz, Tokdemir, and Maraş (Karadeniz et al.,, 2021) | Heart Statlog dataset | Shrunk Covariance classifier | 85.50% | 89.80% | 87.60% | 88.80% |
Anuradha and David (Anuradha and David,, 2021) | Heart Statlog dataset | Majority Voting Ensemble | 87.00% | 87.00% | 87.00% | 87.04% |
Patro, Nayak, and Padhy (Patro et al.,, 2021) | Heart disease UCI dataset | Bayesian Optimized Support Vector Machine | 100.00% | 80.00% | - | 93.30% |
@@
In contrast to the existing literature, mentioned in Table 6, where models were trained on individual heart disease datasets, our approach is novel in that we merged multiple datasets, including the Statlog and UCI datasets, to show the generalizability and robustness of KACQ-DCNN. This is the first attempt in the literature to train a model using such a merged heart disease dataset, offering a more holistic view across different datasets. It’s important to highlight that the Cleveland dataset is a small subset of the UCI dataset (Cleveland, Hungarian, Switzerland, and Long Beach, Virginia), on which several studies trained models individually. Our KAN-enhanced method allows for broader model generalization, potentially capturing patterns across datasets, which could improve robustness in real-world applications. Despite this complexity, our proposed KACQ-DCNN achieved 92% across all evaluation metrics, indicating strong performance even when dealing with the diversity of combined datasets. In the literature, dataset-specific models were used to achieve high accuracy using only one dataset. For example, Mir et al., (2024) got 99.48% accuracy via RF + Adaboost on the Heart Statlog dataset. However, KACQ-DCNN also achieved superior state-of-the-art performance on the Cleveland and Statlog datasets, both sharing an equal maF1 of 96% and the accuracy of 96.23% and 96.30%, respectively.
@@
Over 16 epochs, KACQ-DCNN initially showed consistent improvements in terms of loss and accuracy. The first epoch resulted in a loss of 0.7087, with a validation accuracy of 73.55%. By the third epoch, validation accuracy reached 88.77%, and by the seventh epoch, it peaked at 92.03%. The accuracy subsequently fluctuated without significant improvements. The learning rate was reduced in epoch 12 to enhance convergence from to . However, validation accuracy could not surpass the 92.03% peak, and training was halted at epoch 16 owing to the early stopping mechanism. As shown in Figure 7(a), the training loss steadily decreased across all epochs, indicating effective learning. However, the validation loss plateaued around epoch 7, with minor fluctuations, suggesting that the model was near its generalization capacity. Despite the learning rate reduction at epoch 12, validation loss showed no substantial gains, confirming the onset of overfitting. Figure 7(b) presents the confusion matrix for the KACQ-DCNN model on the test set, offering a granular view of its classification performance. The matrix reveals that 40.22% of all test instances were correctly identified as ’Normal’ (True Negatives), and 51.81% were correctly identified as ‘Heart Failure’ (TPs). These values indicate strong predictive capability for both classes. The model misclassified 4.35% of ‘Normal’ instances as ‘Heart Failure’ (FPs), and 3.62% of ‘Heart Failure’ instances as ‘Normal’ (FNs). The relatively low FN rate is particularly important in a clinical setting, as it signifies fewer missed diagnoses of heart failure. The FP rate, while also low, indicates instances where healthy individuals might be flagged for further investigation.
@@
Figures 8(a) and 8(b) illustrate the performance of KACQ-DCNN via the ROC and precision-recall curves, which are critical metrics for evaluating binary classification tasks. The ROC curve demonstrates a high TP rate against the FP rate, indicating that the model effectively distinguishes between the positive and negative classes. An AUC score of 99.73% on the training set and 94.77% on the test set suggested excellent classification ability, with minimal risk of overfitting. Similarly, the precision-recall curve reveals the model’s ability to maintain high precision and recall rates, achieving values of 99.78% in training and 94.89% in testing. This finding indicates that the model is accurate and robust across both datasets, demonstrating its effectiveness in predicting the target classes.
@@
Key observations from Table 5 regarding overall performance benchmarking include:
1.
The proposed KACQ-DCNN 1-Layered model (Accuracy: 92.03%, ROC-AUC: 94.77%) significantly outperforms all 37 benchmark models, including CML, QNN, and other hybrid variants.
2.
Among classical models, CatBoost (86.43% accuracy) demonstrates the best performance, highlighting its effectiveness with categorical features. MLP classifiers also show strong results.
3.
Pure QNN models generally exhibit substantially lower performance and longer inference times compared to classical and hybrid KACQ-DCNN models, indicating current challenges in their practical application for this task. Performance often degraded with increased quantum layers.
4.
Hybrid models like BiLSTMKANnet and QDenseKANnet (both achieving 88.00% accuracy) show promise, outperforming many CML models and bridging the gap towards the more advanced KACQ-DCNN.
4.1.3 Ablation study analysis